A distribution of values is normal with a mean of 47.1 and a standard deviation of 24.2.

Find P53, which is the score separating the bottom 53% from the top 47%.


P53 =


Enter your answer as a number accurate to 1 decimal place. Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.
Answer #1

P(X < A) = P(Z < (A - mean)/standard deviation)

Mean = 47.1

Standard deviation = 24.2

P(X < P53) = 0.53

P(Z < (P53 - 47.1)/24.2) = 0.53

Take z score corresponding to 0.53 from standard normal distribution table.

(P53 - 47.1)/24.2 = 0.075

P53 = 48.915